OpenStudy (anonymous):

Please help! :( A mathematics journal has accepted 14 articles for publication. However, due to budgetary restraints only 9 articles can be published this month. How many ways can the journal editor assemble 9 of the 14 articles for publication? A. 726,485,760 B. 14 C. 2,002 D. 126

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Thank you, that is what I thought it was as well.

OpenStudy (whpalmer4):

Assuming the problem is asking how many different combinations of articles can the editor make, without caring about the ordering, the answer should be "14 choose 9" or \[\frac{14!}{9!(14-9)!} = \frac{14!}{9!*5!} =\]

OpenStudy (anonymous):

Are you sure it's B?

OpenStudy (whpalmer4):

Let's take a simple case: there are 3 articles: A, B, C and only space to publish 2. The possible combinations are: AB AC BC 3 combinations. Let's test the formula: \[\frac{3!}{2!*1!} = \frac{3*2*1}{2*1*1} = 3\]Checks out. Did I say B? I said you have to evaluate "14 choose 9" and gave the formula...

OpenStudy (whpalmer4):

\[14! = 14*13*12*11*10*9*8*7*6*5*4*3*2*1\]\[9! = 9*8*7*6*5*4*3*2*1\]\[5!=5*4*3*2*1\]Hopefully it is apparent that the 9! in the denominator cancels out much of 14! leaving just \[\frac{14*13*12*11*10}{5*4*3*2*1}=\]

OpenStudy (anonymous):

A

OpenStudy (anonymous):

A

OpenStudy (whpalmer4):

The editor isn't interested in permutations, only combinations.